# Customize-A-Video: One-Shot Motion Customization of Text-to-Video Diffusion Models

### v1

@@ 4 Experiments
Base T2V models
Our methods are applicable to general T2V diffusion models. In the following experiments, we employ the ModelScope T2V model [43] as the pre-trained base model. All videos are pre-processed and generated for 2 seconds, 8 FPS and $256\times 256$ resolution. Other training hyperparameters and model size statistics can be found in the appendix.
Datasets
Since there is no dataset exactly following our single video motion customization setup, we select videos from multiple sources, including LOVEU-TGVE-2023 [49], WebVid-10M [2] and DAVIS [35] datasets to evaluate our method. We also apply our method on in-the-wild videos and demonstrate its generalization ability.
Comparison methods
To our best knowledge there is no previous approach that performs the identical task of one-shot video motion customization with motion variety as we proposed. We pick the most relevant work, Tune-A-Video [48] and Video-P2P [27], which appends raw temporal layers to pre-trained T2I models and finetune both the spatial and temporal attentions on a single reference video prior to subsequent editing. It is worth noting that they additionally rely on DDIM inverted reference video latent as the input during inference and thus only produce temporally deterministic videos with frame layout and fixed view angle. When we remove this additional input they recover some motion generalization while sacrificing temporal coherency. Besides, we compare our method against the pre-trained T2V model, i.e. ModelScope, to prove that our method enhances the base model to produce faithful motions following the reference video that are not trivial to depict via simply prompt engineering.
@@
Method Text Temp. Div.$\uparrow$
Align.$\uparrow$ Consist.$\downarrow$
ModelScope [43] $\underline{31.484}$ $0.175$ $\boldsymbol{0.647}$
Tune-A-Video [48] $31.141$ $0.180$
$\sim$ w/o DDIM Inv $30.304$ $0.206$ $0.348$
Video-P2P [27] $31.001$ $\boldsymbol{0.162}$
$\sim$ w/o DDIM Inv $30.876$ $0.251$ $0.469$
Ours No AA $31.348$ $0.178$ $0.604$
Ours S-LoRA AA $30.816$ $\underline{0.164}$ $0.612$
Ours TextInv AA $\boldsymbol{32.194}$ $0.167$ $0.623$
Ours Both AA $31.449$ $0.171$ $\underline{0.631}$
Quantitative metrics
We measure the performance quantitatively over a subset of videos from [49] which has standard 2-second clips with both ground truth and modification captions. We consider comparisons in terms of three metrics: Text alignment between the generated video frames and the inference prompt, in the form of CLIPScore [15]; Temporal consistency between consecutive frames of the generated video, in the form of LPIPS [53]; Diversity among multiple generated videos with the same prompt and different random noise, in the form of LPIPS. It involves both the appearance and motion diversity by collating aligned frames at the same timestamp.
Figure 4: Ablation results on applying LoRAs on different attention layers. When transferring the camera movement, S-LoRA memorizes the indoor furniture and wall decorations and T-LoRA converts paintings to entrances and sofas to pool benches.
@@ 4.1 Motion Customization from Single Video
Qualitative Results
Fig. 3 illustrates the comparative visual results of one-shot motion customization. The base ModelScope T2V model, although proficient in inferring general motion concepts from its extensive pre-training on large-scale datasets, fails to accurately replicate specific motions guided by reference videos. Contrastingly, Tune-A-Video [48] and Video-P2P [27] leverage DDIM inverted latents extracted from reference videos, resulting in temporally deterministic output with structural constraints by the reference frame layouts. In contrast to both of them, our approach demonstrates the capability to transfer the modeled reference motion to new scenarios and subjects while introducing temporal variations through random noise input. The last row of each example presents two output samples generated by our method, prompted by the same textual cue. Our outcomes not only exhibit diverse subject appearances and background scenes but also showcase variability in motion attributes such as action range, intensity, velocity, and camera perspective.
@@
The quantitative results and comparisons are listed in Tab. 1. In need of universally applicable motion similarity metric encompassing not only human actions but also non-human object and camera movements, text alignment serves to gauge motion accuracy relative to the provided text prompt and inadequately reflects the nature of customized iconic reference motion transfer. Video-P2P leads on temporal consistency with DDIM inverted latent input, but deteriorates considerably in its absence. Ours exhibits comparable coherency against these deterministic methods and the foundation T2V model.
Due to the lack of generic motion representations irrespective of visual appearance, we calculate the per-frame diversity among generated videos that registers both spatial and temporal variety. Comparative evaluation unveils that base ModelScope provides the highest diversity while our methods sacrifice it subtly to gain significant improvements in accurately customizing the exemplar motion, as well as retaining rich varieties in motion details. Tune-A-Video and Video-P2P have deterministic results with DDIM inverted input and yield more yet not as much diverse output without it.
Figure 5: Ablation results on training T-LoRAs with different types of appearance absorbers. No AA remains the bare hands and white sleeves from the original appearance. Both AA reaches the best spatial clearance and displays tidy wall and desk in addition to the evident new dress-up.
@@ 4.2 Ablations Studies
LoRAs on Non-temporal Attentions
While it is intuitive to apply LoRA on only temporal attention layers to learn video motions without original appearance, we also validate the effects of applying LoRA on the spatial attentions only (S-LoRA), or on both spatial and temporal attentions (ST-LoRA) in the base T2V model. Fig. 4 displays the visualization that adding LoRA to spatial attentions significantly impairs the modeling of motions. The models with spatial customization modules primarily memorize the video based on spatial layout, resulting in a substantial degradation of both video appearance and motion customization compared to our T-LoRA approach.
Figure 6: Video Appearance Customization. (1) Reference video. (2) Result of image customization S-LoRA only fails to transfer the motion. (3) Result of motion customization. (4) Result combined both motion and appearance customizations.
Comparisons across Appearance Absorbers
Our method explores four different configurations of Appearance Absorbers (AA). No AA: no appearance absorber is used, S-LoRA AA: a spatial LoRA based appearance absorber is used, TextInv AA: a textual inversion based appearance absorber is used, and Both AA: two appearance absorbers of both above types are used. The comparison results are shown in Fig. 5. No AA remains some original appearance in addition to modeling the motion. S-LoRA AA and TextInv AA are both able to capture the pure action with minimal appearance leakage. We notice that S-LoRA AA is easier to overfit and sometimes causes spatial artifacts while TextInv AA might tend to underfit and leave spatial residues on the other hand. We attribute these properties to the spatial structure of S-LoRA weights inside U-Net blocks while textual inversion works via a 1D learnable embedding as new tokens. Both AA unites their advantages and leads to a comprehensive result with both the reference motion and new appearance clearly reflected.
@@ 5 Applications
With the plug-and-play nature of LoRA, we propose three downstream applications that demonstrate the potential of our models.
5.1 Video Appearance Customization
Our motion customization modules work on the temporal layers and thus can cooperate with image customization methods to manipulate both the temporal and spatial layers in the base T2V model at the same time. In Fig. 6, we inject a Temporal LoRA to present the reference action as well as an image spatial LoRA to reflect the comic style in one comprehensive output. Intermediate results with one LoRA loaded has merely either the spatial or temporal aspect customized on the other hand.
5.2 Multiple Motion Combination
T-LoRA is applied to the original layers with residual connection. Therefore we can customize the base model with multiple T-LoRA modules trained on different reference videos to integrate various movements into one outcome. Fig. 7 demonstrates that our approach can merge the human action of ”falling back” and camera movement of ”dolly zoom” into one target scenario using two T-LoRA modules, achieving combined motions in the generated video.
Figure 7: Multiple Motion Combination. (1-2) two reference videos. (3) Result only customized from the first reference video, only displaying a normal zoom in instead of the dolly zoom, where the foreground remain fixed in scale and the background zoom in or out. (4) Result with two T-LoRAs trained from both reference videos, yielding videos with merged motions. The scale of the humanoid barely changes through the frames while the trees fast zooms out.
5.3 Third-Party Appearance Absorbers
Our staged training pipeline enables loading third-party image customization modules pre-trained on in-the-wild image data as a ready appearance absorber when they share the similar appearance. This skips the first training stage and facilitates reusing modules across videos. In Fig. 8, the reference video focuses on a certain fictional character, and we load a public image LoRA of this character shared in online communities without tuning on the unordered video frames. Trained with its enhancement our T-LoRA avoids the leakage of the original headgear to the replaced subject.
Figure 8: Third-Party Appearance Absorbers. (1) Reference video, whose subject is a popular fictional character. (2) Result trained with no AA. (3) Result trained with a third-party S-LoRA that is pre-trained on in-the-wild images of the character rather than the frames of the reference video.
@@ 6 Conclusion
We introduce the one-shot motion customization task that learns the motion signature from a single reference video and transfer it to new scenes and subjects with variety in both appearance and motion. We propose Temporal LoRA to model the target motion by adding LoRA residual weights on the temporal attention layers of a pre-trained text-to-video diffusion model. We further propose Appearance Absorbers to decouple the spatial information from the reference video so that Temporal LoRA can focus on motion modeling. Extensive experiments demonstrate that our methods yield faithful and diverse videos compared to both per-frame video editing approaches and the base T2V model. Our method is plug-and-play and supports various downstream tasks including precise video editing, video appearance customization and multiple motion combination.
Supplementary Material
Appendix A Implementation Details
A.1 Bypassing Temporal Layers in T2V Models
Many diffusion-based T2V models [3, 14, 43] have their denoising network structure adapted from T2I U-Net with injecting temporal convolution and attention layers. The new temporal layers are usually implemented as residual connections. The models are also usually trained on image and video datasets jointly to acquire both appearance and motion generative capability.
@@ A.1 Bypassing Temporal Layers in T2V Models
A.2 Model Hyperparameters
LoRA [19] typically features very few additional parameters attached to the base model. Its rank $r$ controls the shape of the residual matrix, and $\alpha$ represents its scale when added to the pre-trained model weights. In experiments we discovered that setting the rank of T-LoRA $r_{T}=4$ and the rank of S-LoRA in the appearance absorber $r_{T}=1$ yields satisfactory results. Meanwhile, we empirically determined the alpha values $\alpha_{T}=1$ for T-LoRAs and $\alpha_{S}=0.5$ for S-LoRAs. For textual inversion [10] as the appearance absorber, we set the length of new learnable tokens to 4 and initialize them with 4 words depicting the static content in the frames. Under these settings, our T-LoRA model has merely 0.83M float parameters, compared to the full base T2V model’s 552M dedicated temporal parameters [43].
Figure 9: Comparison between our methods with DDIM inverted reference latent input and Tune-A-Video. Our models are also able to produce precise frame-wise editing output.
Appendix B More Results
B.1 Generation Results
More visual results generated by our models are displayed in Fig. 10. We present two random output samples for each reference video.
B.2 DDIM Inverted Latent Input
Our method can easily incorporate additional deterministic control signals to perform precise video editing. The comparison results to Tuna-A-Video[48] are shown in Fig. 9. Our models prove to be also able to benefit from the DDIM inverted latent of the reference video and yield output stuck to the exact original frame structures.
Figure 10: Additional generation results of our model.
Appendix C Limitations and Future Work
C.1 Spatial Domain Shift
In the first training stage the standalone finetuning of partial layers might have the risk of breaking the consistency among the pre-trained weights if the appearance absorbers are overfitted on static content reconstruction. If the reference frames are out of the pre-trained generalization capacity, the spatial customization might shift the output domain and the subsequent temporal layers will not be able to parse the altered feature maps properly.
C.2 Small Actions and Poses
@@
Some reference motions are intrinsically highly associated with poses, such as walking, running and sitting, and an image can primarily represent them. When the appearance absorbers have modeled the static postures to fit the appearance, in the second training stage the T-LoRA might have little left to learn such as the trivial perturbation across frames. Our model could fail to restore the desired motions evidently under such circumstance.
C.3 Future Plans

### v3

@@ 4 Experiments
Base T2V models.
Our methods are applicable to general T2V diffusion models. In the following experiments, we hire the ModelScope T2V model [45] as the pre-trained base model. All videos are pre-processed and generated for 2 seconds, 8 FPS and $256\times 256$ resolution. Training hyperparameters, model size statistics and time consumption analyses are detailed in the supplementary material.
Datasets.
We select videos from mixed sources, including LOVEU-TGVE-2023 [51], WebVid-10M [2] and DAVIS [36] datasets to evaluate our method. [51] provides ground truth captions and target editing prompts, while we create those for videos from other sources. We also apply our method on in-the-wild videos and demonstrate its generalization ability.
Comparison methods.
As of a new task of one-shot video motion customization, we mainly compare to Tune-A-Video [50] and Video-P2P [28], which append raw temporal layers to pre-trained T2I models and finetune them on a single reference video. It is worth noting that they additionally rely on DDIM inverted reference video latent as the input during inference and thus only produce temporally deterministic videos with fixed frame layout and view angle, so we also evaluate their variants removing this condition. We also compare our method against the pre-trained T2V model, i.e. ModelScope, to prove that our method enhances the base foundation model to produce faithful motions following the reference video that are not trivial to depict via prompt engineering. Besides, we run the concurrent work MotionDirector [60] using their released training code with the same configuration as ours, including the same base T2V model and LoRA hyperparameters for fair comparison.
Quantitative metrics.
We measure the performance quantitatively over a subset containing 53 videos out of [51] of 2-3 seconds with standard original and editing captions. We consider comparisons in terms of three metrics: text alignment between the generated video frames and the inference prompt gauges both generated appearance and motion accuracy, in the form of CLIPScore [15] that associates text and image in a unified space; temporal consistency between consecutive frames of the generated video indicates the generated motion quality, in the form of LPIPS [55] that measures deep feature distance; diversity among multiple generated videos with the same prompt and different random noise involves both spatial and temporal diversity by collating aligned frames at the same timestamp, in the form of LPIPS. It is calculated on 4 random samples per reference video.
User Study.
We conduct a human user study among five algorithms with stochastic output: ModelScope, Tune-A-Video and Video-P2P without DDIM inverted latent, our method with both S-LoRA and textual inversion as the appearance absorbers, and MotionDirector. Every participant is presented 10 random reference videos from [51] and their output videos. Each algorithm outputs two videos, and participants are asked to assess their motion fidelity and motion diversity respectively, from 1 (worst) to 5 (best) stars. Details of the questionnaire design is provided in the supplementary materials.
@@
[…]
ModelScope [45] $31.705$ $0.175$ $\bm{0.636}$
Tune-A-Video [50] $31.149$ $0.185$
$\sim$ w/o DDIM Inversion $30.304$ $0.206$ $0.348$
Video-P2P [28] $31.001$ $\underline{0.162}$
$\sim$ w/o DDIM Inversion $30.876$ $0.251$ $0.469$
MotionDirector [60] $\underline{32.500}$ $0.163$ $0.606$
$\sim$ w/ comparable #params $31.842$ $0.166$ $0.595$
Ours No AA $31.687$ $0.166$ $0.613$
Ours S-LoRA AA $31.913$ $0.163$ $0.618$
Ours TextInv AA $\bm{32.632}$ $\bm{0.160}$ $0.621$
Ours Dual AA $32.193$ $0.164$ $\underline{0.631}$
@@ 4.1 Motion Customization from Single Video
Qualitative Results.
Fig. 3 illustrates the comparative visual results of one-shot motion customization. The base ModelScope T2V model fails to accurately replicate the specific motions as reference guided by simply the text prompt. On the other hand, Tune-A-Video [50] and Video-P2P [28] leverage DDIM inverted latents extracted from reference videos and produce temporally deterministic output with structural constraints by the reference frame layouts. In contrast to both of them, our approach demonstrates the capability to transfer the reference motion to new scenarios and subjects while introducing temporal variations via random noise input. Our outcomes not only exhibit diverse subject appearances and background scenes but also showcase variability in motion attributes such as action range, intensity, velocity, and camera perspective. Concurrent MotionDirector [60] is also able to generate adapted output with variety, while for the hard examples with complex or intensive motion, such as the juggling and firework in Fig. 3, it yields less competitive visual quality than ours. Our T-LoRA is always trained with well-optimized AAs, while MotionDirector might have appearances leaked into the temporal module when their spatial path is jointly being tuned. We also notice that real-world videos usually have divergent spatial and temporal complexity, and our dedicated tuning procedures with independent steps and other schedules for each module reach their individual optimal.
@@
The quantitative results and comparisons are listed in Tab. 1. Our methods outperform the base ModelScope [45], Tune-A-Video [50], Video-P2P [28] and concurrent MotionDirector [60] on both text alignment and temporal consistency. ModelScope [45] provides the highest comprehensive diversity as a foundation model and loses in motion fidelity with text guidance only. Our methods sacrifice it subtly to gain significant improvements in faithfully customizing the exemplar motion, as well as retaining rich varieties in motion details. Tune-A-Video [50] and Video-P2P [28] have the minimal diversity as strictly constrained by frame structures, although at the cost of which they can achieve acceptable temporal consistency. MotionDirector [60] shows comparable text alignment and temporal consistency to ours but its diversity falls behind.
User Study.
We involved 20 evaluators participating our user study and collected 153 and 156 valid ratings per algorithm on each benchmark. The averaged scores are listed in Tab. 2. Our method leads on both motion fidelity and motion diversity. We asked users to rate pure motion fidelity and diversity which are hard to assess by automatic metrics in need of generic motion representations irrespective to spatial structures. These results are complementary to the text alignment and diversity measured in Tab. 1 that mix spatial and temporal signals.
Table 2: Human user study results on [51] dataset. Methods are evaluated from 1 (worst) to 5 (best) stars on each benchmark.
[…]
ModelScope [45] $2.03$ $2.97$
Tune-A-Video [50] w/o DDIM Inversion $2.29$ $2.23$
Video-P2P [28] w/o DDIM Inversion $2.29$ $2.01$
MotionDirector [60] $\underline{3.33}$ $\underline{3.50}$
Ours (w/ dual AA) $\bm{3.72}$ $\bm{3.72}$
Figure 4: Left: Ablations on applying LoRAs on different attention layers. S-LoRA memorizes the indoor furniture and wall decorations and T-LoRA converts paintings to entrances and sofas to pool benches. Right: Ablations on training T-LoRAs with different types of appearance absorbers. No AA adds stylish glasses and the logo but remains most the original appearance. S-LoRA AA and TextInv AA significantly boost the quality while resulting in the strips on the wall and the partially white sleeves. Dual AA reaches best spatial clearance with clear costume and background.
@@ 4.2 Ablations Studies
LoRAs on Non-temporal Attentions.
While it is intuitive to apply LoRA on only temporal attention layers to learn video motions without original appearance, we also validate the effects of applying LoRA on the spatial attentions only (S-LoRA), or on both spatial and temporal attentions (ST-LoRA) in the base T2V model. Fig. 4 left displays the visualizations in which adding LoRA to spatial attentions significantly impairs the motion modeling. The models with spatial elements primarily memorize the video by its spatial layout, resulting in a substantial degradation of both appearance and motion adaptation.
Comparisons among Appearance Absorber Types.
We explore four different configurations of Appearance Absorbers (AA). No AA: no appearance absorber is used, S-LoRA AA: a spatial LoRA based appearance absorber is used, TextInv AA: a textual inversion based appearance absorber is used, and Dual AA: two appearance absorbers of both above types are used. The comparison results are unveiled by Fig. 4. No AA remains some original appearance in addition to modeling the motion. S-LoRA AA and TextInv AA are both able to capture the pure action with minimal appearance leakage. We notice that S-LoRA AA is easier to overfit and sometimes causes spatial artifacts while TextInv AA might tend to underfit and leave spatial residues on the other hand. We attribute these properties to the spatial structure of S-LoRA weights inside U-Net blocks while textual inversion works via a 1D learnable embedding as new tokens. Dual AA unites their advantages and leads to a comprehensive result with both the reference motion and new appearance clearly reflected.
Figure 5: Left: Video appearance customization with both T-LoRA and existing pre-trained S-LoRA (from [43]). Right: Multiple motion combination with two T-LoRAs loaded at the same time. When the robot is slowly jogging, it fast zooms in while the background trees rapidly zoom out (dolly zoom).
@@ 5 Applications
With the plug-and-play nature of LoRA and our staged training pipeline, we present four downstream applications that demonstrate the collaborative potential of our proposed modules.
5.0.1 Video Appearance Customization
Our motion customization module works on temporal layers and thus can cooperate with image customization approaches to manipulate both the temporal and spatial layers in the base T2V model at the same time. In Fig. 5 left, we inject a T-LoRA to present the reference action as well as an image spatial LoRA to reflect the comic style in one comprehensive output.
5.0.2 Multiple Motion Combination
T-LoRA is applied to the original layers with residual connections. Therefore we can customize the base model with multiple T-LoRA modules trained on different source videos to integrate assorted motions into one outcome. Fig. 5 right demonstrates that our method merges the human action of jogging and camera movement of dolly zoom into one target scenario using two T-LoRA modules.
5.0.3 DDIM Inverted Latent Input
Our method can easily incorporate additional deterministic controls to perform precise video editing. The comparison results to Tuna-A-Video[50] are shown as Fig. 6 left. Our models prove to be also able to benefit from the DDIM inverted latent of the reference video and yield output that reproduces the exact original frame structures.
5.0.4 Third-Party Appearance Absorbers
Our staged training pipeline enables reusing appearance absorbers across videos or loading third-party image customization modules as ready appearance absorbers when they share the similar appearance. This skips the first training stage and extends appearance absorber categories to those demand more training data. In Fig. 6 right, we finetune the spatial layers of the UNet following Dreambooth [39] and Eq. 1 on other photos of the same dog. Our T-LoRA trained with it avoids the leakage of the dog color to the white wolf, compared to that trained with the original base model.
Figure 6: Left: Precise frame-wise video editing with DDIM inverted latent as the inference input noise. Right: A third-party Dreambooth UNet pre-tuned on in-the-wild images of the same subject serves as the appearance absorber to train the T-LoRA.
@@ 6 Conclusion
We introduce the one-shot motion customization task that learns the motion signature from a single reference video and transfers it to new scenes and subjects with variety in both appearance and motion. We propose Temporal LoRA to model the target motion by adding LoRA residual weights on the temporal attention layers of a pre-trained text-to-video diffusion model. We further propose Appearance Absorbers to decouple the spatial information from the reference video so that Temporal LoRA can focus on motion modeling. Extensive experiments demonstrate that our methods yield faithful and diverse videos compared to both per-frame video editing approaches and the base T2V model. Our method’s is plug-and-play nature supports various downstream tasks including precise video editing, video appearance customization, multiple motion combination as well as third-party appearance absorbers.
Appendix 0.A Implementation Details
0.A.1 Bypassing Temporal Layers in T2V Models
Many diffusion-based T2V models such as [3, 14, 45, 5] have their denoising network structure adapted from T2I UNet with temporal convolution and attention layers injected. The new temporal layers are usually implemented as residual connections. The models are also usually trained on image and video datasets jointly to acquire both appearance and motion generative capability.
@@ 0.A.1 Bypassing Temporal Layers in T2V Models
0.A.2 Patch Training of Appearance Absorbers
@@
Some motions are intrinsically highly associated with postures, such as walking, running and sitting, and one image can primarily represent them. When the appearance absorbers have modeled the static postures to fit the appearance in the first training stage, T-LoRA might have little left to learn such as only the trivial perturbations across frames.
Therefore, we propose to crop the unordered frames into patches and encourage the appearance absorbers to mainly capture local shapes and textures in the first training stage. This prevent our appearance absorbers from overfitting on the global structures fundamentally. In practice, we find that setting the crop ratio randomly between $0.33$ to $0.67$ yields the best effect to retain the desired motion evidently in the second training stage.
0.A.3 Attentions and LoRAs in T2V Diffusion Models
A base T2V diffusion model involves spatial self-attention (SSA) between a frame and itself, spatial cross-attention (SCA) between each frame and the text prompt, and temporal cross-frame attention (TCFA) among a pixel across all time in each 3D UNet block. We display their computations in Fig. 7. Three types of input and their corresponding $K$, $Q$ and $V$ are marked in respective colors. SSA is calculated between each frame and itself ($K$, $Q$, $V$). SCA is between a frame and the text prompt ($K$, $Q$, $V$). TCFA is among pixels of all frames ($K$, $Q$, $V$). Our LoRAs are applied to all attention weights $W_{*}$ ($\Delta W_{k}$, $\Delta W_{q}$, $\Delta W_{v}$).
Figure 7: (a) Spatial self-attention between each frame and itself; (b) Spatial cross-attention between each frame and the text prompt; (c) Temporal cross-frame attention among pixels of all frames in a video. Batch size is omitted for simplicity.
0.A.4 Model Hyperparameters and Training Time
LoRA [19] typically features very few additional parameters attached to the base model. Its rank $r$ controls the shape of the residual matrix, and $\alpha$ represents its scale when added to the pre-trained model weights. In experiments we discovered that setting the rank of T-LoRA $r_{T}=4$ and the rank of S-LoRA in the appearance absorber $r_{S}=1$ yields satisfactory results. Meanwhile, we empirically determined the alpha values $\alpha_{T}=1$ for T-LoRAs and $\alpha_{S}=0.5$ for S-LoRAs. For textual inversion [10] as the appearance absorber, we set the length of new learnable tokens to 2-6 depending on the content complexity.
Table 3: Quantitative and model size comparison with concurrent work.
Method Text Temp. Div.$\uparrow$ LoRA Rank #Params
Align.$\uparrow$ Consist.$\downarrow$  Temp.  Spat.  Temp.  Spat.
Ours No AA $31.687$ $0.166$ $0.613$ 4 - 831.5K -
Ours S-LoRA AA $31.913$ $0.163$ $0.618$ 1 207.5K
Ours TextInv AA $\bm{32.632}$ $\bm{0.160}$ $\underline{0.621}$ 4K
Ours Both AA $32.193$ $0.164$ $\bm{0.631}$ 1 211.5K
MotionDirector [60] $31.842$ $0.166$ $0.595$ 2 1 779.5K 274.5K
MotionDirector [60] $\underline{32.500}$ $\underline{0.163}$ $0.606$ 4 1 1559K 274.5K
We run experiments on a single NVIDIA RTX A5000 GPU with half-precision floats. Our T-LoRA takes approximately 7 minutes to converge. S-LoRA takes around 0.5 minute and the textual inversion takes 1 minute to converge in the first training stage. This is comparable to Tune-A-Video [50] (6 minutes), Video-P2P [28] (8 minutes in fast mode; 14 minutes in full mode on A6000 for bigger VRAM) and concurrent work MotionDirector [60] (8 minutes) on the same device for the same frame resolution and clip length. The learning rate is set to $5\times 10^{-4}$ for T-LoRA and $5\times 10^{-5}$ for appearance absorbers to prevent overfitting.
It worth noting that due to the difference in LoRA applications between our method and concurrent work MotionDirector, the quantities of parameters and module sizes are not aligned with the same LoRA rank. We apply LoRAs on temporal cross-frame attentions (TCFAs), while MotionDirector moreoever add them to the following feed-forward networks (FFNs). This lead to approximately twice the quantity of parameters to tune. We apply LoRAs on all spatial attentions including the self-attentions (SSAs) and the cross-attentions (SCAs). MotionDirector excludes the SCAs and additionally involves the following FFNs. Thus our spatial LoRAs have comparable amounts of parameters. Tab. 3 expands the quantitative comparison with these model size differences.
Appendix 0.B More Visualizations
0.B.1 Video Generation Results
More video results generated by our models are displayed in Fig. 8. We present two random output samples for each reference video.
Figure 8: Additional generation results of our method.
0.B.2 Appearance Absorber Results
We exhibit the output of our appearance absorbers trained on unorder reference frames with the spatial text prompt in Fig. 9. The 2nd and 4th rows show the generation results with the appearance absorbers (S-LoRA and textual inversion respectively, same below) loaded and all temporal layers bypassed in the base T2V model, and the spatial part of the text prompt is used. It yields individual static frame replicating the reference appearances with random postures. The dynamic information is successfully left for our temporal customization module to learn in the next stage. The 3rd and 5th rows show the output videos with the appearance absorbers loaded on the full base T2V model, and the full text prompt is used. With the temporal description, the model can still only produce generic motions upon the learned appearances, indicating the necessity and effectiveness of our temporal customization module training. It can be further noticed that S-LoRA and textual inversion have different flavors of spatial modeling due to their different mechanisms, and thus loading both of them achieves the best performance with comprehensive and thorough appearance absorbing.
Figure 9: Appearance absorbers’ generation output. The 2nd and 3rd rows have S-LoRA loaded. The 4th and 5th rows have textual inversion loaded. The training prompts and special tokens are noted above for each sample.
0.B.3 Training Schedule Variances
Our modules fit on each reference video individually to model its unique motion signal. Fig. 10 displays cases whose optimal iterations vary across different reference videos. In general we observed that the convergence steps increase along with the complexity the specific reference motion and that of the original appearance.
We also observed that different types of appearance absorbers may exhibit different characteristics that affect the optimal checkpoint step and the output details. In Fig. 11 we present some cases where appearance absorbers vary in their effect of assisting following stages to capture the accurate motion or to generate novel scenes in certain iterations.
Figure 10: Examples where different reference videos require different tuning iterations. (1-2) Simpler motions such as camera movements usually converge faster. (3) More complicated motions such as animal or human actions would demand more tuning steps.
Figure 11: Examples where different appearance absorbers exhibit different characteristics. (1) Our Both AA absorbs the original appearance more thoroughly, leading to more diverse new background generated. (2) Our Both AA may reduce the necessary convergence step compared to a single appearance absorber. (3) Our Both AA may be more stable and enable more tuning iterations without collapse to thoroughly clean up the original art style and generate a new one.
Appendix 0.C User Study Questionnaire Design
We present example questions in our user studies in Fig. 12. Every reference video is presented with the output videos by random 4 out of 5 algorithms to be evaluated. For motion fidelity, 1-star represents the most dissimilar and 5-star represents the most faithful transferred motions w.r.t. the reference video. For motion diversity, 1-star indicates the most identical and 5-star indicates the most diverse generated motions among the two output videos.
Figure 12: An example question in the human user study. Participants are asked to rate each algorithm’s output videos from 1 to 5 stars.
Appendix 0.D Limitation Discussions
Per Instance Finetuning.
Our method tunes on each reference video individually. Similar to comparing approaches [60, 50, 28], our method needs specialized recipes for different videos of diverse appearances and motions. The training configurations and iterations depend on the target video and can vary a lot as analysed in Sec. 0.B.3. The trade-off balance between the object motion fidelity and its diversity also relies on dedicated hyperparameters and adjustments between underfitting and overfitting, like its image customization counterparts [39, 10] have described. Though, our staged training pipeline and plug-and-play designs enable reusing both the appearance absorbers and T-LoRAs for future training and compositional inference, which improve their usability.
Spatial Domain Shift.
The standalone finetuning of partial layers might have the risk of breaking the consistency among the pre-trained weights if the appearance absorbers overfit on static content reconstruction. If the reference frames are out of the T2V model’s pre-trained generalization capacity, the spatial customization might shift its output domains during training and the subsequent temporal layers will be unable to parse the altered feature maps properly in the next stage. We suggest smaller learning rate and LoRA scale to pick the checkpoint when the reference video has complex appearances such as uncommon contents or extraordinary styles. Applying our methods on advanced base T2V models with leading capabilities also helps.
Text Encoding Conflict.
While extensive spatial customization modules can be alternatively utilized as our appearance absorbers, some of them might encounter text mapping conflict when collaborating with the temporal customization modules. For example, we choose not to apply LoRA on the text encoder in T2V diffusion models although it can enhance the spatial modelling and appearance decomposition. Modifications on the pre-trained text encoder could tamper the original mapping from text to its embedding, and then T-LoRA will learn the motion associated with the altered text tokens. Finally it might not be triggered properly by the vanilla text encoder without the appearance absorbers during inference. The null-text prompt training trick for LoRA without triggering words might help to handle this issue.
Appendix 0.E Future Work